find a queadraric model for each set of values
(-1,6) (1,4) (2,9)

Answer by scott8148(6628)

y = ax^2 + bx + c

6 = a - b + c

4 = a + b + c

9 = 4a + 2b + c

solve the system of equations for the coefficients a, b, and c